If z is a complex number and |z + 16| = 4 |z + l|, find the value of | z|


1
Expert's answer
2021-07-22T15:11:58-0400

z=x+yi


|z+16|=4|z+1|


|x+yi+16|²=4²|x+yi+1|²


[(x+16)²+y2]=16[(x+1)²+y²]


x²+32x+256+y²=16x²+32x+16+16y²


15x²+15y²=240


x²+y²=16


|z|=|x+yi|=4